# Efficient Shapley Value-based Non-Uniform Pruning of Large Language Models

### v2

@@ 4.2 Shapley Value Approximation
Figure 3: Selection of Transformer coalitions under an LLM with 8 Transformers.
To address the NP-hard problem, this subsection proposes a sliding window-based Shapley value (SWSV) approximation method to estimate the contribution of Transformers. The key is that many coalitions of Transformers are useless and meaningless when evaluating LLMs. We consider an LLM with 8 Transformers as shown in Figure 3, where Transformers 1 and 8 are active and Transformers 2-7 are masked. We have two limitations: 1) too few active Transformers; and 2) too far apart between active Transformers.
Specifically, a well-trained LLM relies on maintaining a sufficiently large network structure, even in the presence of sparsity, to preserve its representation capacity and functionality Li et al. (2022); Lee et al. (2024). Sparsity, while effective in reducing computational costs, must be implemented judiciously to ensure that the structural integrity and critical pathways of the model remain intact. An LLM with an insufficient number of active Transformer layers—such as configurations involving only two active layers (e.g., Transformers 1 and 8)—is highly likely to suffer from significant performance degradation. This occurs because such configurations disrupt the hierarchical processing of information, resulting in incomplete or suboptimal feature representations. Moreover, when many intermediate Transformers are masked, the direct transmission of outputs from an earlier Transformer (e.g., Transformer 1) to a much later one (e.g., Transformer 8) bypasses the essential intermediate processing steps. This shortcut diminishes the model’s ability to learn and propagate nuanced information through the intermediate layers, leading to a distorted representation of the Shapley value. Such scenarios are problematic as they fail to reflect the contributions of individual layers within the model architecture.
To address these issues, it is critical to design an SV approximation method to avoid extreme patterns with few Transformers. The goal is to improve the accuracy of SV estimation, enhance model performance, and reduce computational complexity by preserving critical Transformers.
Figure 4: Illustration of the proposed approximation.
To this end, we propose a sliding window-based Shapley value approximation method for evaluating the contribution of Transformers in LLMs (Figure 4). When calculating the SV of Transformer $t\in\mathcal{T}$, we consider the sliding window with $N=|\mathcal{S}_{t}|$ Transformers, where these Transformers are closely connected. Figure 4 shows the sliding window with 3 Transformers, where the SV of Transformer 1 is only related to the Transformers in this sliding window. It is expressed as:
@@
Neighboring Transformers around a target Transformer $t$ play a significant role in determining its SV, as they strongly influence the context and interactions that contribute to the target Transformer’s functionality. To effectively capture these interdependencies, a sliding window strategy is employed, which selects a set of neighboring Transformers both before and after the target Transformer. This approach ensures that the SV calculation takes into account the most relevant local interactions within the Transformer stack. The sliding window size $N$ is generally set to an odd number, such as 3, 5, or 7, to maintain symmetry around the target Transformer. For instance, with a sliding window of size 3, the calculation includes one preceding Transformer ($t-1$), the target Transformer itself ($t$), and one succeeding Transformer ($t+1$). This balanced selection ensures that contributions from both upstream and downstream Transformers are equally considered, facilitating a more accurate evaluation of the target Transformer’s importance.
Input: $\mathcal{T}$, $N$, an empty set $\mathcal{V}$, and a dataset;
Output: Shapley values of Transformers;
1
2while $t\in\mathcal{T}$ do
@@ 17      Calculate $\hat{\phi}_{t}$ by (2);
Special considerations are made for boundary Transformers, such as the first and last ones, where the default sliding window size might result in insufficient neighbors. In these cases, the sliding window is adjusted to include the available Transformers, while maintaining the specified number of neighbors as much as possible. For example, when evaluating the first Transformer, additional downstream Transformers can be included to compensate for the absence of preceding Transformers. Similarly, for the last Transformer, upstream Transformers are prioritized. This ensures consistency in the Shapley value calculations across all Transformers, preserving the intended neighborhood size while respecting the structural boundaries of the model. The pseudocode of SWSV is shown in Algorithm 1. Its computational complexity is $\mathcal{O}(T2^{N-1})$, which is much lower than the original complexity.
4.3 SWSV for Pruning LLMs
After calculating the contributions of Transformers (${\hat{\phi}_{t}},{t\in\mathcal{T}}$), these contributions can be utilized to allocate the pruning ratio for each Transformer. Transformers with greater contributions are considered more important and thus have smaller pruning ratios. Therefore, the pruning ratio $\rho_{t}$ is inversely proportional to the contribution, i.e., $\rho_{t}\propto 1/\hat{\phi}_{t}$. However, since overly sparse Transformers can degrade the performance of LLMs, it is necessary to constrain the allocated pruning ratios. To achieve this, we introduce a small positive value $\lambda$ (e.g., 0.05 or 0.1) as a boundary. The allocated pruning ratios are restricted to lie within the given pruning ratio range, with an allowable error of $\lambda$, i.e., $\rho_{t}\in[\rho-\lambda,\rho+\lambda]$. The detailed mathematical expression is:
$\rho_{t}=\rho-a_{t}+\text{mean}\{a_{t}\}_{t\in\mathcal{T}},$ (3)
SV-NUP is a framework which can be applied to advanced pruning methods (e.g., Wanda Sun et al. (2023) and SparseGPT Frantar and Alistarh (2023)). The pseudocode of SV-NUP is presented in Algorithm 2.
@@ 4.3 SWSV for Pruning LLMs
1 Calculate the contributions $\{\hat{\phi}_{t}\}_{t\in\mathcal{T}}$ of Transformers by SWSV;
@@
4       Prune the Transformer $t$ by an advanced pruning method at the pruning ratio $\rho_{t}$;
@@ 5 Experimental Evalution
5.1 Experiment Settings
We evaluate the proposed SV-NUP method on the LLaMA-V1, LLaMA-V2, and OPT model families Frantar and Alistarh (2023). Due to computational resource constraints, our focus is on the 7-billion and 13-billion parameter variants. All LLMs are quantized with float16 and the sequence length is set as 2048. To assess the performance of the pruned LLMs, we employ two general metrics: PPL and zero-shot evaluation Sun et al. (2023). PPL reflects the model’s ability to predict the next word given the preceding context Muhlgay et al. (2023), where lower values indicate better performance. We evaluate PPL using the WikiText-2 dataset Merity et al. (2016), selecting two randomly downloaded subsets of WikiText-2 for testing. For zero-shot evaluation, we choose 4 tasks—BoolQ, RTE, WinoGrande (WG), and OpenBookQA (OBQA)—from the EleutherAI LM Harness Mihaylov et al. (2018), where higher scores indicate better performance.
Table 1: PPL↓  performance of LLaMA family pruned by 3 existing methods with/without (w./wo.) SV-NUP at 70% sparsity. We highlight the improved performance in blue.
LLaMA-7B LLaMA-13B LLaMA2-7B LLaMA2-13B
@@ 5.1 Experiment Settings
SV-NUP w. SparseGPT 15.10 (-18.01%) 11.06 (-19.55%) 14.84 (-17.00%) 12.48 (-13.27%)
5.2 Comparison Baselines
We select 3 baselines for comparison: 1) Magnitude Jaiswal et al. (2024a), 2) Wanda Sun et al. (2023), and 3) SparseGPT Frantar and Alistarh (2023). The proposed SV-NUP method is integrated into these baselines, allowing us to directly observe the performance improvement brought by it. In addition, we compare SV-NUP with OWL Yin et al. (2023), a pruning ratio allocation framework based on outliers.
We implement all experiments in PyTroch and use the HuggingFace library to download the pre-trained LLMs and datasets. All experiments are performed on a server with 4 NVIDIA A100 GPUs, 1 AMD EPYC CPU, and 252 GB of memory. The limitation $\lambda$ of the pruning ratio is set to 0.1. The size of sliding windows $N$ is set as 3, 5, or 7.
5.3 PPL Improvement by SV-NUP
@@
We first evaluate the PPL performance gains achieved by integrating SV-NUP into LLM pruning methods, as shown in Table 1. SV-NUP significantly enhances existing pruning approaches, particularly Magnitude and Wanda, while SparseGPT demonstrates minimal improvements due to its already strong baseline performance. SV-NUP has 61.28 % improvement on LLaMA-13B by Wanda. Wanda consistently benefits from SV-NUP across all LLaMA models, with the largest impact observed in larger models like LLaMA-13B, where SV-NUP effectively mitigates performance degradation. SparseGPT, despite its high baseline effectiveness, achieves further refinement with SV-NUP, highlighting the compatibility and synergy between the two. Overall, SV-NUP emerges as a crucial enhancement, stabilizing pruning outcomes and delivering superior results, especially when combined with advanced methods like SparseGPT.
@@
Table 2: PPL↓  performance of LLaMA family and OPT family pruned by different methods at 50% sparsity. We highlight the best performance in bold.
TinyLLaMA LLaMA-7B Vicuna-7B LLaMA2-7B LLaMA3.2-3B OPT-2.7B OPT-6.7B
Dense 7.7735 5.6772 6.9031 5.4721 7.8137 12.4705 10.8602
Magnitude 21.6367 17.2882 24.0034 16.0301 139.4124 265.2033 968.7209
OWL w. Magnitude 18.7673 16.3453 21.3578 15.7421 99.2163 207.0598 363.6955
SV-NUP w. Magnitude 18.1645 15.9755 20.7409 15.3144 89.4710 158.2674 308.4234
Wanda 10.8549 7.0884 8.5325 6.7741 12.6977 13.9611 12.0780
OWL w. Wanda 11.1533 7.0941 8.5731 6.7954 12.6735 14.7622 12.3921
SV-NUP w. Wanda 11.0834 7.0610 8.4428 6.7959 12.1578 14.0323 12.0065
SparseGPT 9.8810 6.8701 8.0767 6.6000 11.0357 12.0591 11.2711
OWL w. SparseGPT 10.1521 6.8785 8.2247 6.6388 10.9828 13.3490 11.5101
SV-NUP w. SparseGPT 10.1308 6.8336 8.0762 6.5843 10.8495 12.9768 11.2582
@@
Table 3: PPL↓  and classification performance of Vicuna-7B and OPT-6.7B pruned by different methods at 50% sparsity. The ”mean” is the average of BoolQ, RTE, WG, and OBQA.
PPL↓ BoolQ RTE WG OBQA Mean
Vicuna-7B Dense 6.9031 0.7810 0.6823 0.6938 0.3460 0.6258
Magnitude 24.0034 0.5462 0.5271 0.5817 0.2200 0.4687
OWL w. Magnitude 21.3578 0.6067 0.5784 0.6056 0.2300 0.5052
SV-NUP w. Magnitude 20.7409 0.6037 0.5884 0.5959 0.2360 0.5060
Wanda 8.5325 0.6969 0.6859 0.6519 0.2960 0.5827
OWL w. Wanda 8.5731 0.6315 0.7040 0.6685 0.2960 0.5750
SV-NUP w. Wanda 8.4428 0.6976 0.6968 0.6630 0.2880 0.5863
SparseGPT 8.0767 0.6437 0.6534 0.6480 0.2940 0.5598
OWL w. SparseGPT 8.2247 0.6780 0.6715 0.6472 0.2840 0.5702
SV-NUP w. SparseGPT 8.0762 0.6945 0.6823 0.6567 0.3020 0.5839
OPT-6.7B Dense 10.8602 0.6606 0.5523 0.6519 0.2760 0.5352
Magnitude 968.7209 0.3804 0.5271 0.5059 0.1760 0.3974
OWL w. Magnitude 363.6955 0.3813 0.5271 0.5564 0.2140 0.4197
SV-NUP w. Magnitude 308.4234 0.3908 0.5271 0.5567 0.2120 0.4217
Wanda 12.0780 0.6214 0.5271 0.6014 0.2380 0.4970
OWL w. Wanda 12.3921 0.6220 0.5271 0.5864 0.2380 0.4934
SV-NUP w. Wanda 12.0065 0.6211 0.5271 0.6069 0.2420 0.4993
SparseGPT 11.2711 0.6358 0.5343 0.6433 0.2540 0.5168
OWL w. SparseGPT 11.5101 0.6324 0.5307 0.6393 0.2320 0.5086
SV-NUP w. SparseGPT 11.2582 0.6410 0.5543 0.6338 0.2540 0.5208
Table 4: PPL↓  performance under different $N$ at 50% sparsity.
@@
SV-NUP w. Magnitude 16.0491 16.0081 15.9755
SV-NUP w. Wanda 7.09322 7.06104 7.09215
SV-NUP w. SparseGPT 6.8382 6.83358 6.84875
@@
SV-NUP w. Magnitude 525.407 308.423 661.884
SV-NUP w. Wanda 12.0469 12.0339 12.0065
SV-NUP w. SparseGPT 11.2766 11.2661 11.2582
Table 5: Pruning ratio allocation by two SV approximation methods on LLaMA-7B and OPT-6.7B at 50% sparsity. The ”mean” is the average of BoolQ, RTE, WG, and OBQA.
PPL↓ BoolQ RTE WG OBQA Mean
@@
Figure 5: Distribution of the allocated pruning ratio on LLaMA-7B by 3 methods at 50% sparsity.
5.4 PPL and Zero-shot Performance by Baselines
The results in Table 2 underscore the effectiveness of integrating SV-NUP and OWL into existing pruning methods for the LLaMA and OPT model families at 50% sparsity. Across most scenarios, SV-NUP consistently demonstrates superior performance. Specifically, Magnitude combined with SV-NUP consistently outperforms both its standalone version and Magnitude with OWL. Notably, while SV-NUP slightly improves the performance of Wanda and SparseGPT, the gains are less pronounced, as the pruned LLMs at 50% sparsity already exhibit robust performance, leaving limited room for further enhancements. Interestingly, SV-NUP’s effectiveness is more apparent with larger LLMs, such as the 7-billion-parameter models, whereas its impact diminishes for smaller models like TinyLLaMA and OPT-2.7B. For Wanda, the integration of SV-NUP results in a notable reduction in PPL compared to Wanda alone, highlighting the synergy between SV-NUP and fine-grained pruning strategies. OWL also contributes positively, particularly when integrated with SparseGPT and Wanda, though its impact is slightly less consistent than that of SV-NUP. Overall, integrating SV-NUP stabilizes and enhances performance across various pruning methods, emerging as a particularly strong addition. It consistently ensures competitive results and maintains lower PPL across a wide range of LLM architectures.
Table 3 presents the PPL and classification performance of OPT-2.7B and OPT-6.7B models pruned by different methods at 50% sparsity. For OPT-2.7B, SV-NUP with SparseGPT achieves a mean score of 0.4951, which is slightly lower than SparseGPT alone (0.5039), suggesting that the integration of SV-NUP with SparseGPT may not provide performance gains and could slightly degrade results for smaller OPT models. In contrast, for OPT-6.7B, SV-NUP with SparseGPT achieves a mean score of 0.5208, which is comparable to SparseGPT alone (0.5168), indicating that SV-NUP can maintain performance levels when combined with SparseGPT for larger models.
Table 4 illustrates the impact of varying sliding window sizes ($N$) on performance. The results indicate that $N=5$ and $N=7$ achieve lower PPL for both LLaMA-7B and OPT-6.7B. However, the choice of $N$ is highly sensitive to different LLMs and pruning methods, highlighting the importance of selecting an appropriate window size tailored to specific scenarios.
5.5 Different SV Approximation Methods
We compare our proposed SWSV method with the existing SV approximation method proposed in Kolpaczki et al. (2024) to demonstrate its efficiency, as shown in Table 5. For LLaMA-7B, SWSV with SparseGPT achieves the highest mean score (0.5584), slightly surpassing SV with SparseGPT (0.5493), showcasing its ability to enhance performance. Similarly, for OPT-6.7B, SWSV with SparseGPT achieves a mean score of 0.5208, outperforming SV with SparseGPT (0.5111). These results demonstrate that SWSV not only provides marginal performance improvements over standard SV, particularly when combined with SparseGPT, but also offers lower computational complexity. This highlights its potential as a more efficient and effective sparsity allocation approach.
Figure 5 shows the distribution of allocated pruning ratios for LLaMA-7B at 50% sparsity across three methods: SWSV, SV, and OWL. SWSV demonstrates significant variability, particularly in earlier layers, with pruning ratios fluctuating more widely before stabilizing closer to the latter layers.

### v3

@@ 4.2 Shapley Value Approximation
Figure 4: Selection of Transformer coalitions under an LLM with 8 Transformer layers.
To address the NP-hard problem, this subsection proposes a SWSV method to estimate the contribution of Transformer layers. The key is that many coalitions of Transformer layers are useless and meaningless when evaluating LLMs. We consider an LLM with 8 Transformer layers as shown in Figure 4, where Transformers 1 and 8 are active and Transformer layers 2-7 are masked. We have two limitations: 1) too few active Transformer layers; and 2) too far apart between active Transformer layers.
Specifically, a well-trained LLM relies on maintaining a sufficiently large network structure, even in the presence of sparsity, to preserve its representation capacity and functionality [34; 35]. Sparsity, while effective in reducing computational costs, must be implemented judiciously to ensure that the structural integrity and critical pathways of the model remain intact. An LLM with an insufficient number of active Transformer layers—such as only two active layers (e.g., Transformer layers 1 and 8)—is highly likely to suffer from significant performance degradation. This occurs because such configurations disrupt the hierarchical processing of information, resulting in incomplete or suboptimal feature representations. Moreover, when many intermediate Transformer layers are masked, the direct transmission of outputs from an earlier Transformer layer (e.g., Transformer layer 1) to a much later one (e.g., Transformer layer 8) bypasses the essential intermediate processing steps. This shortcut diminishes the model’s ability to learn and propagate nuanced information through the intermediate layers, leading to a distorted representation of the Shapley value. Such scenarios are problematic as they fail to reflect the contributions of individual layers within the model architecture.
To address these issues, it is critical to design an SV approximation method to avoid extreme patterns with few Transformer layers. The goal is to improve the accuracy of SV estimation, enhance model performance, and reduce computational complexity by preserving critical Transformer layers.
Figure 5: Illustration of SWSV.
To this end, we propose a sliding window-based Shapley value approximation method for evaluating the contribution of Transformer layers in LLMs (Figure 5). When calculating the SV of Transformer layer $t\in\mathcal{T}$, we consider the sliding window with $N=|\mathcal{S}_{t}|$ Transformer layers, where these Transformer layers are closely connected. Figure 5 shows the sliding window with 3 Transformer layers, where the SV of Transformer layer 1 is only related to the Transformer layers in this sliding window. It is expressed as:
@@
Input: $\mathcal{T}$, $N$, $\mathcal{V}$, and a dataset;
Output: Shapley values $\{\hat{\phi}_{t}\}_{t\in\mathcal{T}}$;
1 while $t\in\mathcal{T}$ do
@@ 16      Calculate $\hat{\phi}_{t}$ by (2);
Neighboring Transformer layers around a target Transformer layer $t$ play a significant role in determining its SV, as they strongly influence the context and interactions that contribute to the target Transformer layer’s functionality. To effectively capture these interdependencies, a sliding window strategy is employed, which selects a set of neighboring Transformer layers both before and after the target Transformer layer. This approach ensures that the SV calculation takes into account the most relevant local interactions within the Transformer layer stack. The sliding window size $N$ is generally set to an odd number, such as 3, 5, or 7, to maintain symmetry around the target Transformer layer. For instance, with a sliding window of size 3, the calculation includes one preceding Transformer layer ($t-1$), the target Transformer layer itself ($t$), and one succeeding Transformer layer ($t+1$). This balanced selection ensures that contributions from both upstream and downstream Transformer layers are equally considered, facilitating a more accurate evaluation of the target Transformer layer’s importance.
@@
1 Calculate the contributions $\{\hat{\phi}_{t}\}_{t\in\mathcal{T}}$ of Transformer layers by SWSV;
@@
4       Prune the Transformer layer $t$ by an advanced pruning method at the pruning ratio $\rho_{t}$;
@@
Special considerations are made for boundary Transformer layers, such as the first and last ones, where the default sliding window size might result in insufficient neighbors. In these cases, the sliding window is adjusted to include the available Transformer layers, while maintaining the specified number of neighbors as much as possible. For example, when evaluating the first Transformer layer, additional downstream Transformer layers can be included to compensate for the absence of preceding Transformer layers. Similarly, for the last Transformer layer, upstream Transformer layers are prioritized. This ensures consistency in the Shapley value calculations across all Transformer layers, preserving the intended neighborhood size while respecting the structural boundaries of the model. The pseudocode of SWSV is shown in Algorithm 1. Its computational complexity is $\mathcal{O}(T2^{N-1})$, significantly lower than the original $\mathcal{O}(2^{T})$. For instance, computing contributions for LLaMA-7B with $N=5$ requires only 512 subsets, compared to $4,294,967,296$ in the original approach.
4.3 SV-NUP for LLMs
After calculating the contributions of Transformer layers (${\hat{\phi}_{t}},{t\in\mathcal{T}}$), these contributions can be utilized to allocate the pruning ratio for each Transformer layer. Transformer layers with greater contributions are considered more important and thus have smaller pruning ratios. Therefore, the pruning ratio $\rho_{t}$ is inversely proportional to the contribution, i.e., $\rho_{t}\propto 1/\hat{\phi}_{t}$. However, since overly sparse Transformer layers can degrade the performance of LLMs, it is necessary to constrain the allocated pruning ratios. To achieve this, we introduce a small positive value $\lambda$ (e.g., 0.05 or 0.1) as a boundary. The allocated pruning ratios are restricted to lie within the given pruning ratio range, with an allowable error of $\lambda$, i.e., $\rho_{t}\in[\rho-\lambda,\rho+\lambda]$. The detailed mathematical expression is:
$\rho_{t}=\rho-a_{t}+\text{mean}\{a_{t}\}_{t\in\mathcal{T}},$ (3)
SV-NUP is a framework which can be applied to advanced pruning methods (e.g., Wanda [14] and SparseGPT [4]). The pseudocode of SV-NUP is presented in Algorithm 2.
@@ 5 Experimental Evalution
LLMs and Baselines. We evaluate SV-NUP on the LLaMA-V1/v2/v3, and OPT model families [4]. Due to computational constraints, we select LLMs with up to 13 billion parameters, quantized to float16 with a sequence length of 2048. We select 3 baselines for comparison: 1) Magnitude (2017) [36], 2) Wanda (2024) [14], and 3) SparseGPT (2023) [4]. The proposed SV-NUP method can be integrated into these baselines, allowing us to directly assess its performance gains. In addition, we compare SV-NUP with OWL (2024) [12] and ALS (2024) [37], two pruning ratio allocation frameworks based on outliers and information orthogonality, respectively. We implement all experiments in PyTorch and use the HuggingFace library to download the pre-trained LLMs and datasets. All experiments are conducted on a server equipped with 4 NVIDIA A100 GPUs (40GB each), one AMD EPYC CPU, and 252 GB of memory. The pruning ratio constraint $\lambda$ is set to 0.1, and the sliding window size $N$ is chosen as 3, 5, or 7. Notably, since ALS does not publicly release its source code, we report its results directly from the original paper.
Experiment Settings. We implement all experiments in PyTorch and use the HuggingFace library to download the pre-trained LLMs and datasets. All experiments are performed on a server with 4 NVIDIA A100 GPUs, 1 AMD EPYC CPU, and 252 GB of memory. To assess the performance of the pruned LLMs, we employ two general metrics: PPL and zero-shot evaluation [14]. PPL reflects the model’s ability to predict the next word given the preceding context [38], where lower values indicate better performance. We evaluate PPL using the WikiText-2 dataset [39], selecting two randomly downloaded subsets of WikiText-2 for testing. For zero-shot evaluation, we choose 4 tasks—BoolQ, RTE, WinoGrande (WG), and OpenBookQA (OBQA)—from the EleutherAI LM Harness [40], where higher scores indicate better performance. Detailed settings are shown in Appendix B.
Table 1: PPL↓  performance of LLaMA-v1/v2 family pruned by 3 existing methods with (w.) SV-NUP at 70% sparsity. We highlight the improved performance in blue.
Model LLaMA-7B LLaMA-13B LLaMA2-7B LLaMA2-13B
@@
SparseGPT w. SV-NUP 15.10 (-18.01%) 11.06 (-19.55%) 14.84 (-18.01%) 12.48 (-18.01%)
@@
PPL Improvement by SV-NUP. We first evaluate the PPL performance gains achieved by integrating SV-NUP into LLM pruning methods, as shown in Table 1. SV-NUP significantly enhances existing pruning approaches, particularly Magnitude and Wanda, while SparseGPT demonstrates minimal improvements due to its already strong baseline performance. SV-NUP has 61.28 % improvement on LLaMA-13B by Wanda. Wanda consistently benefits from SV-NUP across all LLaMA models, with the largest impact observed in larger models like LLaMA-13B, where SV-NUP effectively mitigates performance degradation. SparseGPT, despite its high baseline effectiveness, achieves further refinement with SV-NUP, highlighting the compatibility and synergy between the two. Overall, SV-NUP emerges as a crucial enhancement, stabilizing pruning outcomes and delivering superior results, especially when combined with advanced methods like SparseGPT.
PPL and Zero-shot Performance. The results in Table 2 underscore the effectiveness of integrating SV-NUP, OWL, and ALS into existing pruning methods for the LLaMA and OPT model families at 50% sparsity. Particularly, as ALS does not release its source code, the results of ALS are not fair and are for reference only. In most scenarios, SV-NUP consistently demonstrates superior performance. Specifically, Magnitude combined with SV-NUP consistently outperforms both its standalone version and Magnitude with OWL. Notably, while SV-NUP slightly improves the performance of Wanda and SparseGPT, the gains are less pronounced, as the pruned LLMs at 50% sparsity already exhibit robust performance, leaving limited room for further enhancements. Interestingly, SV-NUP’s effectiveness is more apparent with larger LLMs, such as the 7-billion-parameter models, whereas its impact diminishes for smaller OPT-2.7B.
@@
Table 2: PPL↓  performance of LLaMA-v1/v2/v3 family and OPT family pruned by different methods at 50% sparsity. We highlight the best performance in bold. We set the results of ALS as gray.
Model LLaMA-7B Vicuna-7B LLaMA2-7B LLaMA3.2-3B OPT-2.7B OPT-6.7B
Dense 5.6772 6.9031 5.4721 7.8137 12.4705 10.8602
Magnitude 17.2882 24.0034 16.0301 139.4124 265.2033 968.7209
Magnitude w. ALS 16.8000 15.1900 950.0000
Magnitude w. OWL 16.3453 21.3578 15.7421 99.2163 207.0598 363.6955
Magnitude w. SV-NUP 15.9755 20.7409 15.3144 89.4710 158.2674 308.4234
Wanda 7.0884 8.5325 6.7741 12.6977 13.9611 12.0780
Wanda w. ALS 12.4700 11.6100 19.1600
Wanda w. OWL 7.0941 8.5731 6.7954 12.6735 14.7622 12.3921
Wanda w. SV-NUP 7.0610 8.4428 6.7959 12.1578 14.0323 12.0065
SparseGPT 6.8701 8.0767 6.6000 11.0357 12.0591 11.2711
SparseGPT w. ALS 11.8700 10.9900 12.2900
SparseGPT w. OWL 6.8785 8.2247 6.6388 10.9828 13.3490 11.5101
SparseGPT w. SV-NUP 6.8336 8.0762 6.5843 10.8495 12.9768 11.2582
Table 3 compares PPL of pruning methods (Magnitude, Wanda, SparseGPT) combined with ALS, OWL, and SV-NUP across LLaMA-7B, LLaMA2-7B, and OPT-6.7B at 30–50% sparsity. SV-NUP achieves the lowest PPL in nearly all cases, demonstrating consistent superiority. For example, on OPT-6.7B at 50% sparsity, SV-NUP attains a PPL of 308.4234 (Magnitude) and 12.0065 (Wanda), markedly better than OWL (363.6955, 12.3921) and ALS (950.0000, 19.1600). The gains are especially pronounced at higher sparsities and for larger models, highlighting SV-NUP’s robustness and scalability. These results solidify SV-NUP as a state-of-the-art pruning technique for efficient model compression.
Table 3: PPL↓  performance of LLaMA-7B, LLaMA2-7B, and OPT-6.7B at different sparsities.
Model LLaMA-7B LLaMA2-7B OPT-6.7B
Sparsity 30% 40% 50% 30% 40% 50% 30% 40% 50%
Magnitude w. ALS 16.8000 9.6000 11.0300 15.1900 950.0000
Magnitude w. OWL 6.7413 8.8245 16.3453 6.3403 8.1432 15.7421 12.8504 20.4357 363.6955
Magnitude w. SV-NUP 6.6769 8.5035 15.9755 6.3438 8.2723 15.3144 12.3753 18.7027 308.4234
Wanda w. ALS 12.4700 9.1500 9.8100 11.6100 19.1600
Wanda w. OWL 6.0066 6.3583 7.0941 5.7712 6.0827 6.7954 10.7365 11.2408 12.3921
Wanda w. SV-NUP 5.9885 6.3294 7.0610 5.7554 6.0702 6.7959 10.6790 11.1051 12.0065
SparseGPT w. ALS 11.8700 9.1100 9.6700 10.9900 12.2900
SparseGPT w. OWL 5.9512 6.2543 6.8785 5.7599 6.0340 6.6388 10.9449 11.0805 11.5101
SparseGPT w. SV-NUP 5.9483 6.2285 6.8336 5.7535 6.0306 6.6243 10.8853 10.9279 11.2582
@@
Table 4: Zero-shot performance↑  of LLaMA-7B and LLaMA2-7B pruned by different methods at 50% sparsity. The "mean" is the average of BoolQ, RTE, WG, and OBQA.
Model LLaMA-7B LLaMA2-7B
Zero-shot task BoolQ RTE WG OBQA Mean↑ BoolQ RTE WG OBQA Mean↑
Dense 75.08% 66.06% 70.09% 34.20% 61.36% 77.71% 62.82% 69.14% 31.40% 60.27%
Magnitude 54.56% 54.15% 59.43% 22.60% 47.68% 62.94% 57.04% 63.38% 26.80% 52.54%
Magnitude w. ALS 59.82% 54.51% 61.25% 36.60% 53.05% 71.38% 55.24% 65.19% 41.40% 58.30%
Magnitude w. OWL 58.69% 55.23% 61.56% 27.80% 50.82% 63.88% 53.79% 62.90% 28.60% 52.29%
Magnitude w. SV-NUP 57.63% 57.76% 61.40% 27.20% 51.00% 64.92% 55.96% 64.64% 28.60% 53.53%
Wanda 72.45% 55.40% 66.61% 28.60% 55.77% 75.81% 54.87% 68.35% 31.20% 57.56%
Wanda w. ALS 73.70% 60.65% 66.30% 38.60% 59.81% 75.47% 54.87% 67.80% 44.80% 60.74%
Wanda w. OWL 73.43% 53.43% 65.98% 30.60% 55.86% 77.22% 54.51% 68.51% 30.40% 57.66%
Wanda w. SV-NUP 72.91% 55.23% 67.56% 29.80% 56.38% 76.18% 54.15% 69.40% 31.20% 57.73%
SparseGPT 74.43% 49.82% 68.11% 28.20% 55.14% 70.46% 54.51% 67.01% 28.00% 55.00%
SparseGPT w. ALS 74.28% 54.87% 66.77% 39.00% 58.73% 70.98% 55.96% 67.96% 40.00% 58.73%
SparseGPT w. OWL 72.84% 54.51% 67.88% 26.00% 55.31% 71.50% 56.32% 68.19% 28.40% 56.10%
SparseGPT w. SV-NUP 73.30% 54.51% 68.75% 26.80% 55.84% 68.84% 63.54% 67.96% 28.60% 57.23%
Table 4 presents the zero-shot performance of various pruning methods applied to LLaMA-7B and LLaMA2-7B across four downstream tasks. The unpruned dense models serve as baselines, achieving 61.36% and 60.27% average accuracy, respectively. Among the pruning strategies, magnitude pruning yields the most significant performance degradation, whereas both Wanda and SparseGPT show stronger resilience, with SparseGPT with SV-NUP achieving up to 57.23% on LLaMA2-7B, closely matching the dense baseline. Notably, our proposed method SV-NUP consistently improves performance across all pruning strategies and model backbones. It is worth noting that the results of ALS occasionally outperform the dense models; we suspect this is due to differences in experimental settings or hyperparameters rather than true pruning benefits. Thus, ALS results are reported for reference only and are excluded from direct comparison. Overall, the results validate the effectiveness and robustness of SV-NUP in mitigating performance loss introduced by non-uniform pruning.
Table 5: PPL↓  under different $N$ at 50% sparsity.
@@
Magnitude w. SV-NUP 16.0491 16.0081 15.9755
Wanda w. SV-NUP 7.0932 7.0610 7.0921
SparseGPT w. SV-NUP 6.8382 6.8336 6.8487
@@
Magnitude w. SV-NUP 525.4074 308.4234 661.8835
Wanda w. SV-NUP 12.0469 12.0339 12.0065
SparseGPT w. SV-NUP 11.2766 11.2661 11.2582
Windows Size Selection. Table 5 reports PPL of pruned LLaMA-7B and OPT-6.7B using different pruning strategies with SV-NUP, evaluated under varying window sizes ($N{=}3,5,7$). Across both models and all pruning methods, the PPL remains largely stable as the window size changes, indicating that SV-NUP is robust to the choice of window size. These results suggest that SV-NUP does not require careful tuning of the window size.
Table 6: Pruning ratio allocation by two SV approximation methods on LLaMA-7B and OPT-6.7B at 50% sparsity.
PPL↓ BoolQ RTE WG OBQA Mean↑
@@
Different SV Approximation Methods. Table 6 shows our proposed SWSV method with the existing SV approximation method in [41] to demonstrate its efficiency. For LLaMA-7B, SWSV with SparseGPT achieves the highest mean score (0.5584), slightly surpassing SV with SparseGPT (0.5493), showcasing its ability to enhance performance. These results demonstrate that SWSV not only provides marginal performance improvements over standard SV, but also offers lower computational complexity.
Figure 6: Activation cosine similarity.
Activation Cosine Similarity. Figure 6 illustrates the activation cosine similarity across Transformer layers between the original LLaMA-7B and its pruned counterparts using Wanda and Wanda w. SV-NUP at 70% sparsity. Notably, Wanda w. SV-NUP consistently yields higher similarity scores than Wanda alone across all layers (0–30), suggesting superior preservation of the original LLaMA-7B activations post-pruning. As pruning error increases, cosine similarity correspondingly declines. Further insights into SV-NUP’s advantages over other baselines are provided in Figure 7 in Appendix C. These findings substantiate the efficacy of SV-NUP in mitigating architectural distortion during pruning. Additional experimental results are presented in Appendix C.